OpenStudy (anonymous):

Please locate the inflection points and relative extrema of y=x-sinx on the interval [0,4pi] using the second derivative test.

OpenStudy (campbell_st):

find the 1st and second derivatives \[\frac{dy}{dx} = 1 - \cos(x)\] \[\frac{d^2y}{dx^2} = \sin(x)\] set \[\frac{dy}{dx} = 0\] to find stationary points then cos(x) = 1 x = 0, 2 pi, 4pi set \[\frac{d^2y}{dx^2} = 0\] sin(x) = 0 so x = 0, pi, 2pi, 3pi, 4pi. test the stationary points from the 1st derivative in the 2nd derivative sin(0) = 0 sin(2pi) = 0 sin(4pi) = 0 then the points where x = 0, 2pi, 4pi are horizontal points of inflexion. test either side of the points for a change in the sign of the 2nd derivative  then the points of inflexion are \[(0,0)**,...(\pi, \pi)...(2\pi, 2\pi)**, .... (3\pi, 3\pi),.... (4\pi, 4\pi)**\] and ** are horizontal points of inflexion